OpenStudy (anonymous):

If you deposit $600 in an account that pays 6.75% compounded continuously, how long will it take to double your investment? (t is time in years) the formula is A=Pe^rt A= amont t=time in yrs. P=principal and r=annual interest SHOW WORK PLEASE =)

OpenStudy (amistre64):

ln(2) ---- perhaps? 1+r

OpenStudy (amistre64):

nah, its just r on the bottom methinks

OpenStudy (amistre64):

\begin{array}{r||} 2P=P.e^{rt}\\ \frac{2P}{P}=e^{rt}\\ ln(2) = rt\\ \frac{ln(2)}{r}=t \end{array}

OpenStudy (anonymous):

amistre64, you're getting lazy

OpenStudy (amistre64):

\begin{align} \color{#ff0000}{a}\color{#0ff000}{s}\ \color{#000ff0}{c}\color{#0000ff}{a}\color{#5f00f0}{n}\text{ see; it aint about the amount at all :)} \end{align}

OpenStudy (amistre64):

lol .... lazy? me? it takes effort to write in this latex stuff lol

OpenStudy (amistre64):

ln(scalar) ------- = how much time it takes r